# A Survey of Large Language Models

### v15

@@ 2.1 Background for LLMs
$\bullet$ Step-by-step reasoning. For small language models, it is usually difficult to solve complex tasks that involve multiple reasoning steps, e.g., mathematical word problems. In contrast, with the chain-of-thought (CoT) prompting strategy [33], LLMs can solve such tasks by utilizing the prompting mechanism that involves intermediate reasoning steps for deriving the final answer. This ability is speculated to be potentially obtained by training on code [47, 33]. An empirical study [33] has shown that CoT prompting can bring performance gains (on arithmetic reasoning benchmarks) when applied to PaLM and LaMDA variants with a model size larger than 60B, while its advantage over the standard prompting becomes more evident when the model size exceeds 100B. Furthermore, the performance improvement with CoT prompting seems to be also varied for different tasks, e.g., GSM8K $>$ MAWPS $>$ SWAMP for PaLM [33].
@@
PanGu-$\Sigma$ [119] Mar-2023 1085 PanGu-$\alpha$ 329B tokens - 512 Ascend 910 100 d $\checkmark$ -
@@ 3.1 Publicly Available Model Checkpoints or APIs
$\bullet$ Mistral. The Mistral series [137, 138] consist of Mistral (7B), Mistral NeMo (12B), Mistral Large 2 (123B), and Mixtral (8$\times$7B and 8$\times$22B), which have been widely known for their strong performance on various mainstream benchmarks (e.g., MMLU and GSM8k). Mistral NeMo is featured with a long context window of 128K at the parameter scale of 12B. Although Mistral NeMo is trained with quantization awareness, it enables FP8 inference without sacrificing performance. Mistral Large 2 is the largest and most powerful model of the Mistral series, which supports 11 natural languages and more than 80 programming languages. Mixtral is a kind of sparse Mixture-of-Experts (SMoE) model that activates only part of the parameters during inference, making it more efficient compared to dense models of the same size.
@@ 3.2 Commonly Used Corpora for Pre-training
[…]
@@ 4.1.3 Data Scheduling
$\bullet$ Mathematics. Llemma [258] is proposed to enhance the mathematical capacities of general-purpose LLMs. It is developed based on CodeLLaMA. Although CodeLLaMA [254] mainly focuses on the coding ability, experiments have shown that it performs better than its base model LLaMA 2 on mathematics benchmarks [258]. Based on CodeLLaMA, Llemma is continually trained on mixtures of scientific papers, web data containing mathematical text and code (2T general tokens $\rightarrow$ 500B code-heavy tokens $\rightarrow$ 50$\sim$200B math-heavy tokens). Note that the pre-training data of Llemma also contains 5% general domain data as a form of regularization.
@@ 4.2.1 Typical Architectures
TABLE VI: Comparison of parallelism and complexity of different models. $T$ represents sequence length, $H$ represents the dimension of the input representation, $N$ represents the dimension after compression in SSMs, and $M$ represents the number of layers in each Hyena module.
@@ 4.2.2 Detailed Configuration
[…]
@@
$\bullet$ FlashAttention. Different from most existing approximate attention methods that trade-off model quality to improve the computing efficiency, FlashAttention [302] proposes to optimize the speed and memory consumption of attention modules on GPUs from an IO-aware perspective. There exist different levels of memory on modern GPUs, e.g., SRAM with a fast IO and HBM with a relatively slow IO. FlashAttention organizes the input into blocks and introduces necessary recomputation, both to make better use of the fast memory SRAM. Implemented as a fused kernel in CUDA, FlashAttention has been integrated into PyTorch [211], DeepSpeed [74], and Megatron-LM [75]. The updated version FlashAttention-2 [303] further optimizes the work partitioning of GPU thread blocks and warps, leading to around 2$\times$ speedup when compared to the original FlashAttention.
@@ 4.2.3 Pre-training Tasks
$\mathcal{L}_{DAE}(\mathbf{x})=\log P(\tilde{\mathbf{x}}|\mathbf{x}_{\backslash\tilde{\mathbf{x}}}).$ (7)
@@ 4.2.5 Summary and Discussion
PanGu-$\alpha$ (200B) - $2\times 10^{-5}$ Adam - 0.1 - -
@@
ERNIE 3.0 Titan (260B) - $1\times 10^{-4}$ Adam FP16 0.1 1.0 -
@@ 5.1.3 The Effect of Instruction Tuning
Models Dataset Mixtures Instruction Numbers Lexical Diversity Chat QA
@@
w/ complexity 70,000 70.43 76.96 39.73 33.25
w/ diversity 70,000 75.59 81.55 38.01 30.03
w/ difficulty 70,000 73.48 79.15 32.55 31.25
w/ scaling 220,000 57.78 51.13 33.81 26.63
@@
w/ complexity 70,000 70.43 77.94 46.89 35.75
w/ diversity 70,000 75.59 78.92 44.97 36.40
w/ difficulty 70,000 73.48 80.45 43.15 34.59
w/ scaling 220,000 57.78 58.12 38.07 27.28
@@ 6 Utilization
In-context Learning (ICL) KATE [420] Demonstration selection (similar; k-NN)
@@
Chain-of-thought Prompting (CoT) Complex CoT [426] Demonstration (complexity-based selection)
@@ 6.1.1 Prompt Creation
D7. As a symbol sequence can typically be divided into multiple segments (e.g., $i_{1},i_{2},i_{3}$longrightarrow $i_{1},i_{2}$ and $i_{2},i_{3}$), the preceding ones can be used as in-context exemplars to guide LLMs to predict the subsequent ones, meanwhile providing historical information. ②③
@@
$<$Demonstrations$>$
Articles: $<$insert articles, each delimited by triple quotes$>$
Question: $<$insert question$>$
@@ 5. What are the missing references mentioned by the individual reviews?
The review texts are below: $<$insert three comments $R_{1}$, $R_{2}$, $R_{3}$ from the reviewers$>$
Meta-review: $<$insert meta-review$>$
$<$Demonstrations$>$
@@
The review texts are below: $<$insert three comments $R_{1}$, $R_{2}$, $R_{3}$ from the reviewers$>$
@@
$<$Demonstrations$>$
Question: $<$insert question$>$
@@ 6.2.1 ICL Formulation
Formally, let $D_{k}=\{f(x_{1},y_{1}),\dots,f(x_{k},y_{k})\}$ represent a set of demonstrations with $k$ examples, where $f(x_{k},y_{k})$ is the prompt function that transforms the $k$-th task example into natural language prompts. Given the task description $I$, demonstration $D_{k}$, and a new input query $x_{k+1}$, the prediction of the output $\hat{y}_{k+1}$ generated from LLMs can be formulated as follows373737 When ICL was introduced in the GPT-3’s paper [55], it was originally defined to be a combination of the task description and demonstration examples, wherein either component is dispensable. Following this definition, when a LLM is required to solve an unseen task by using only task descriptions, it can be also considered to perform ICL for task solving, whereas the ICL ability can be enhanced by instruction tuning. :
@@ 6.4.1 The Overall Framework
In this paradigm, there are typically three components: task planner, plan executor, and environment383838 Despite the similarity with RL, our formulation decouples the planning and execution phases, whereas in RL, they are typically interleaved in the agent. This paradigm is defined in a general yet slightly loose way, and it mainly aims to help readers understand the key idea underlying the planning approaches of LLMs. . Specifically, task planner, which is played by LLMs, aims to generate the whole plan to solve a target task. The plan can be presented in various forms, e.g., an action sequence in the form of natural language [432] or an executable program written in programming language [436]. The LLM-based task planner can be enhanced with the memory mechanism for plan storage and retrieval, which is helpful for long-horizon tasks. Then, plan executor is responsible for executing the actions in the plan. It can be implemented by models like LLMs for textual tasks [434] or by tools like code interpreters for coding tasks [443]. Furthermore, environment refers to where the plan executor carries out the actions, which can be set differently according to specific tasks, e.g., the LLM itself [529] or an external virtual world like Minecraft [530]. It provides feedback about the execution result of the action to the task planner, either in the form of natural language [443] or from other multimodal signals [439].
@@ 7.1.1 Language Generation
Code Synthesis. In addition to generating high-quality natural language text, existing LLMs also show strong abilities to generate formal language, especially computer programs (i.e., code) that satisfy specific conditions, called code synthesis [637]. Unlike natural language generation, as the generated code can be directly checked by execution with corresponding compilers or interpreters, existing work mostly evaluates the quality of the generated code from LLMs by calculating the pass rate against the test cases, i.e., pass@$k$ [^39] Given $k$ programs generated by the LLM, pass@$k$ is computed as 1 when at least one program passes all test cases, or else 0. Recently, several code benchmarks focusing on functional correctness are proposed to assess the code synthesis abilities of LLMs, such as APPS [376], HumanEval [105], and MBPP [223]. Typically, they consist of diverse programming problems, with text specification and test cases for correctness checking. To improve such an ability, it is key to fine-tuning (or pre-training) LLMs on code data, which can effectively adapt LLMs to code synthesis tasks [86]. In addition, existing work has proposed new strategies to generate code, e.g., sampling multiple candidate solutions [223] and planning-guided decoding [638], which can be considered as the imitation of bug-fixing and code-planning processes by programmers. Impressively, LLMs have recently shown competitive performance with humans by achieving a ranking of the top 28% among users on the programming contest platform Codeforces [114]. Further, GitHub Copilot has been released to assist programming in coding IDEs (e.g., Visual Studio and JetBrains IDEs), which can support a variety of languages including Python, JavaScript, and Java. A viewpoint article entitled “The End of Programming” [639] in Communications of the ACM has discussed the impact of AI programming in the field of computer science, emphasizing an important shift towards the highly adaptive LLM as a new atomic unit of computation.
@@ 7.1.2 Knowledge Utilization
Open-Book QA. Unlike closed-book QA, in open-book QA tasks, LLMs can extract useful evidence from the external knowledge base or document collections, and then answer the question based on the extracted evidence [655, 656, 657, 658]. Typical open-book QA datasets (e.g., Natural Questions [556], OpenBookQA [568], and SQuAD [571]) have overlap with closed-book QA datasets, but they incorporate external data sources, e.g., Wikipedia. The metrics of accuracy and F1 score are widely used in open-book QA tasks for evaluation. To select relevant knowledge from external resources, LLMs are often paired with a text retriever (or even a search engine), which is trained independently or jointly with LLMs [655, 659, 81]. Also, previous work [660, 661, 662] has indicated that retrievers can assist LLMs in verifying and rectifying the reasoning path. In evaluation, existing studies mainly focus on testing how LLMs utilize the extracted knowledge to answer the question and show that the retrieved evidence can largely improve the accuracy of the generated answers, even enabling a smaller LLM to outperform 10×10$\times$10 × larger ones [655, 659]. Further, open-book QA tasks can be also employed to evaluate the recency of knowledge information. Pre-training or retrieving from outdated knowledge resources may cause LLMs to generate incorrect answers for time-sensitive questions [655].
@@ 7.1.3 Complex Reasoning
$\bullet$ Numerical computation. For complex reasoning tasks, LLMs still face difficulties in the involved numerical computation, especially for the symbols that are seldom encountered during pre-training, such as arithmetic with large numbers [679, 49, 692]. To tackle this issue, a direct way is to tune LLMs on synthesized arithmetic problems [693, 359]. Also, a surge of studies improve the numerical computation performance by tracing intermediate calculation steps in training and inference stages [678, 694, 359], e.g., scratchpad tracing. In addition, existing work [80] has also incorporated external tools (e.g., calculator), especially for handling arithmetic operations. More recently, ChatGPT has provided a plugin mechanism to use external tools [667]. In this way, LLMs need to learn how to properly manipulate the tools. For this purpose, researchers have augmented the examples using tools (even the LLM itself) for tuning the LLM [695, 80], or devised instructions and exemplars for in-context learning [436]. In addition to the aid of external tools, recent studies find that tokenizing digits into individual tokens (e.g., LLaMA and Galactica tokenizers) is a useful approach to enhancing the inherent arithmetic ability of LLMs [692, 359]. One possible explanation is that subword tokenization techniques can result in inconsistent sequences when tokenizing numbers. For instance, with a subword tokenizer the integer 7481 may be tokenized as $7\_481$, while 74815 may be tokenized as $748\_15$ (the same numerical substrings with different splits) [359]. As a comparison, digit-based tokenization for numbers can avoid such an inconsistency, thus likely improving the numerical computation ability of LLMs.
@@ 9.1.2 Adapting Context Window
$\bullet$ $\Lambda$-shaped context window. Some prior work has revealed that LLMs tend to allocate greater attention weights to the starting and nearest tokens among all previous tokens [947, 948], and it potentially results in the “lost in the middle” phenomenon [949]. Based on this observation, LM-Infinite [950] and StreamingLLM [948] propose to employ a “$\Lambda$-shaped” attention mask, which selectively preserves the initial tokens and the nearest tokens that each query can attend to and then discards any tokens beyond this scope. Experiments demonstrate that this method can facilitate extra-long text generation with a fixed memory [948]. However, it may struggle to model the long-range dependency in the context window, since it cannot effectively utilize the information from the discarded tokens [948].
@@ 9.2 LLM-empowered Agent
Components. Typically, there are three main components in an LLM-based agent: memory, planning474747 Section 6.4 introduces planning as a utilization approach for LLMs, while in this section, we describe its utilization as a functional component in LLM-based agents. , and execution. Specifically, the memory component aims to store the information perceived from the environment and can be utilized to support decision-making. In particular, LLM-based agents usually maintain information in both short-term memory and long-term memory with the operations of reading and writing. Short-term memory usually refers to the internal context window of LLMs (i.e., input), where LLMs can read and write through actions like reasoning [963]. While long-term memory can be mapped to the external storage like vector databases [539], where LLMs can read through retrieval and write with reflection [688]. Specially, profiles are usually implemented with long-term memory, which is an important feature for an agent that specifies its role and function [821]. The planning component is responsible for generating the action plan based on the information from the memory component. In data format, the plan usually takes the form of text-based instructions [434] or code-based programs [436]. To generate it, LLM-based agents will first propose several candidates and then select a more suitable one among them [429]. The initial plan can be further refined with execution feedback from the environment [530]. The execution component is in charge of carrying out the plan from the planning component, which can be fulfilled by the internal LLM [434] or external tools [963].

### v16

@@ 2.1 Background for LLMs
$\bullet$ Step-by-step reasoning. For small language models, it is usually difficult to solve complex tasks that involve multiple reasoning steps, e.g., mathematical word problems. In contrast, with the chain-of-thought (CoT) prompting strategy [33], LLMs can solve such tasks by utilizing the prompting mechanism that involves intermediate reasoning steps for deriving the final answer. This ability is speculated to be potentially obtained by training on code [47, 33]. An empirical study [33] has shown that CoT prompting can bring performance gains (on arithmetic reasoning benchmarks) when applied to PaLM and LaMDA variants with a model size larger than 60B, while its advantage over the standard prompting becomes more evident when the model size exceeds 100B. Furthermore, the performance improvement with CoT prompting seems to be also varied for different tasks, e.g., GSM8K >> MAWPS >> SWAMP for PaLM [33].
@@
PanGu-$\Sigma$ [119] Mar-2023 1085 PanGu-$\alpha$ - - 329B tokens - 512 Ascend 910 100 d $\checkmark$ -
@@ 3.1 Publicly Available Model Checkpoints or APIs
[…]
@@ 3.2 Commonly Used Corpora for Pre-training
 $\bullet\textit{ Book Data.}$ BookCorpus [158] is a commonly used dataset in previous small-scale models (e.g., GPT [122] and GPT-2 [26]), consisting of over 11,000 books covering a wide range of topics and genres (e.g., novels and biographies). Another large-scale book corpus is Project Gutenberg [159], consisting of over 70,000 literary books including novels, essays, poetry, drama, history, science, philosophy, and other types of works in the public domain. It is currently one of the largest open-source book collections, which is used in training of MT-NLG [113] and LLaMA [57]. As for Books1 [55] and Books2 [55] used in GPT-3 [55], they are much larger than BookCorpus but have not been publicly released so far.
@@ 4.1.3 Data Scheduling
[…]
@@ 4.2.1 Typical Architectures
TABLE VI: Comparison of parallelism and complexity of different models. $T$ represents sequence length, $H$ represents the dimension of the input representation, NN represents the dimension after compression in SSMs, and $M$ represents the number of layers in each Hyena module.
@@ 4.2.2 Detailed Configuration
GeLU [279] $\mathrm{GeLU}(\mathbf{x})=\mathrm{0.5}\mathbf{x}\otimes[1+\mathrm{erf}(\mathbf{x}/\sqrt{2})],\text{~~~}\mathrm{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}dt$
@@
[…]
@@ 4.2.3 Pre-training Tasks
[…]
@@ 4.2.5 Summary and Discussion
PanGu-$\alpha$ (200B) - $2\times 10^{-5}$ - - Adam - 0.1 - -
@@
ERNIE 3.0 Titan (260B) - $1\times 10^{-4}$ - - Adam FP16 0.1 1.0 -
@@ 5.1.3 The Effect of Instruction Tuning
Models
Dataset
Mixtures
Instruction
Numbers
Lexical
Diversity
Chat QA
@@
w/ complexity
70,000 70.43 76.96 39.73 33.25
w/ diversity
70,000 75.59 81.55 38.01 30.03
w/ difficulty
70,000 73.48 79.15 32.55 31.25
w/ scaling
220,000 57.78 51.13 33.81 26.63
@@
w/ complexity
70,000 70.43 77.94 46.89 35.75
w/ diversity
70,000 75.59 78.92 44.97 36.40
w/ difficulty
70,000 73.48 80.45 43.15 34.59
w/ scaling
220,000 57.78 58.12 38.07 27.28
@@ 6 Utilization
In-context
Learning (ICL)
KATE [420] Demonstration selection (similar; k-NN)
@@
Chain-of-thought
Prompting (CoT)
Complex CoT [426] Demonstration (complexity-based selection)
@@ 6.1.1 Prompt Creation
D7. As a symbol sequence can typically be divided into multiple segments (e.g., $i_{1},i_{2},i_{3}$ $\longrightarrow$ $i_{1},i_{2}$ and $i_{2},i_{3}$), the preceding ones can be used as in-context exemplars to guide LLMs to predict the subsequent ones, meanwhile providing historical information. ②③
@@
<<Demonstrations>>
Articles: <<insert articles, each delimited by triple quotes>>
Question: <<insert question>>
@@ 5. What are the missing references mentioned by the individual reviews?
The review texts are below: <<insert three comments $R_{1}$, $R_{2}$, $R_{3}$ from the reviewers>>
Meta-review: <<insert meta-review>>
<<Demonstrations>>
@@
The review texts are below: <<insert three comments $R_{1}$, $R_{2}$, $R_{3}$ from the reviewers>>
@@
<<Demonstrations>>
Question: <<insert question>>
@@ 6.2.1 ICL Formulation
Formally, let $D_{k}=\{f(x_{1},y_{1}),\dots,f(x_{k},y_{k})\}$ represent a set of demonstrations with $k$ examples, where $f(x_{k},y_{k})$ is the prompt function that transforms the $k$-th task example into natural language prompts. Given the task description $I$, demonstration $D_{k}$, and a new input query $x_{k+1}$, the prediction of the output $\hat{y}_{k+1}$ generated from LLMs can be formulated as follows [^37] When ICL was introduced in the GPT-3’s paper [55], it was originally defined to be a combination of the task description and demonstration examples, wherein either component is dispensable. Following this definition, when a LLM is required to solve an unseen task by using only task descriptions, it can be also considered to perform ICL for task solving, whereas the ICL ability can be enhanced by instruction tuning. :
@@ 6.4.1 The Overall Framework
In this paradigm, there are typically three components: task planner, plan executor, and environment [^38] Despite the similarity with RL, our formulation decouples the planning and execution phases, whereas in RL, they are typically interleaved in the agent. This paradigm is defined in a general yet slightly loose way, and it mainly aims to help readers understand the key idea underlying the planning approaches of LLMs. . Specifically, task planner, which is played by LLMs, aims to generate the whole plan to solve a target task. The plan can be presented in various forms, e.g., an action sequence in the form of natural language [432] or an executable program written in programming language [436]. The LLM-based task planner can be enhanced with the memory mechanism for plan storage and retrieval, which is helpful for long-horizon tasks. Then, plan executor is responsible for executing the actions in the plan. It can be implemented by models like LLMs for textual tasks [434] or by tools like code interpreters for coding tasks [443]. Furthermore, environment refers to where the plan executor carries out the actions, which can be set differently according to specific tasks, e.g., the LLM itself [529] or an external virtual world like Minecraft [530]. It provides feedback about the execution result of the action to the task planner, either in the form of natural language [443] or from other multimodal signals [439].
@@ 7.1.1 Language Generation
Code Synthesis. In addition to generating high-quality natural language text, existing LLMs also show strong abilities to generate formal language, especially computer programs (i.e., code) that satisfy specific conditions, called code synthesis [637]. Unlike natural language generation, as the generated code can be directly checked by execution with corresponding compilers or interpreters, existing work mostly evaluates the quality of the generated code from LLMs by calculating the pass rate against the test cases, i.e., pass@kk [^39] Given $k$ programs generated by the LLM, pass@$k$ is computed as 1 when at least one program passes all test cases, or else 0. Recently, several code benchmarks focusing on functional correctness are proposed to assess the code synthesis abilities of LLMs, such as APPS [376], HumanEval [105], and MBPP [223]. Typically, they consist of diverse programming problems, with text specification and test cases for correctness checking. To improve such an ability, it is key to fine-tuning (or pre-training) LLMs on code data, which can effectively adapt LLMs to code synthesis tasks [86]. In addition, existing work has proposed new strategies to generate code, e.g., sampling multiple candidate solutions [223] and planning-guided decoding [638], which can be considered as the imitation of bug-fixing and code-planning processes by programmers. Impressively, LLMs have recently shown competitive performance with humans by achieving a ranking of the top 28% among users on the programming contest platform Codeforces [114]. Further, GitHub Copilot has been released to assist programming in coding IDEs (e.g., Visual Studio and JetBrains IDEs), which can support a variety of languages including Python, JavaScript, and Java. A viewpoint article entitled “The End of Programming” [639] in Communications of the ACM has discussed the impact of AI programming in the field of computer science, emphasizing an important shift towards the highly adaptive LLM as a new atomic unit of computation.
@@ 7.1.2 Knowledge Utilization
Open-Book QA. Unlike closed-book QA, in open-book QA tasks, LLMs can extract useful evidence from the external knowledge base or document collections, and then answer the question based on the extracted evidence [655, 656, 657, 658]. Typical open-book QA datasets (e.g., Natural Questions [556], OpenBookQA [568], and SQuAD [571]) have overlap with closed-book QA datasets, but they incorporate external data sources, e.g., Wikipedia. The metrics of accuracy and F1 score are widely used in open-book QA tasks for evaluation. To select relevant knowledge from external resources, LLMs are often paired with a text retriever (or even a search engine), which is trained independently or jointly with LLMs [655, 659, 81]. Also, previous work [660, 661, 662] has indicated that retrievers can assist LLMs in verifying and rectifying the reasoning path. In evaluation, existing studies mainly focus on testing how LLMs utilize the extracted knowledge to answer the question and show that the retrieved evidence can largely improve the accuracy of the generated answers, even enabling a smaller LLM to outperform $10\times$ larger ones [655, 659]. Further, open-book QA tasks can be also employed to evaluate the recency of knowledge information. Pre-training or retrieving from outdated knowledge resources may cause LLMs to generate incorrect answers for time-sensitive questions [655].
@@ 7.1.3 Complex Reasoning
$\bullet$ Numerical computation. For complex reasoning tasks, LLMs still face difficulties in the involved numerical computation, especially for the symbols that are seldom encountered during pre-training, such as arithmetic with large numbers [679, 49, 692]. To tackle this issue, a direct way is to tune LLMs on synthesized arithmetic problems [693, 359]. Also, a surge of studies improve the numerical computation performance by tracing intermediate calculation steps in training and inference stages [678, 694, 359], e.g., scratchpad tracing. In addition, existing work [80] has also incorporated external tools (e.g., calculator), especially for handling arithmetic operations. More recently, ChatGPT has provided a plugin mechanism to use external tools [667]. In this way, LLMs need to learn how to properly manipulate the tools. For this purpose, researchers have augmented the examples using tools (even the LLM itself) for tuning the LLM [695, 80], or devised instructions and exemplars for in-context learning [436]. In addition to the aid of external tools, recent studies find that tokenizing digits into individual tokens (e.g., LLaMA and Galactica tokenizers) is a useful approach to enhancing the inherent arithmetic ability of LLMs [692, 359]. One possible explanation is that subword tokenization techniques can result in inconsistent sequences when tokenizing numbers. For instance, with a subword tokenizer the integer 7481 may be tokenized as 7_4817\_481, while 74815 may be tokenized as 748_15748\_15 (the same numerical substrings with different splits) [359]. As a comparison, digit-based tokenization for numbers can avoid such an inconsistency, thus likely improving the numerical computation ability of LLMs.
@@ 9.1.2 Adapting Context Window
$\bullet$ $\Lambda$-shaped context window. Some prior work has revealed that LLMs tend to allocate greater attention weights to the starting and nearest tokens among all previous tokens [947, 948], and it potentially results in the “lost in the middle” phenomenon [949]. Based on this observation, LM-Infinite [950] and StreamingLLM [948] propose to employ a $\Lambda$-shaped” attention mask, which selectively preserves the initial tokens and the nearest tokens that each query can attend to and then discards any tokens beyond this scope. Experiments demonstrate that this method can facilitate extra-long text generation with a fixed memory [948]. However, it may struggle to model the long-range dependency in the context window, since it cannot effectively utilize the information from the discarded tokens [948].
@@ 9.2 LLM-empowered Agent
Components. Typically, there are three main components in an LLM-based agent: memory, planning [^47] Section 6.4 introduces planning as a utilization approach for LLMs, while in this section, we describe its utilization as a functional component in LLM-based agents. , and execution. Specifically, the memory component aims to store the information perceived from the environment and can be utilized to support decision-making. In particular, LLM-based agents usually maintain information in both short-term memory and long-term memory with the operations of reading and writing. Short-term memory usually refers to the internal context window of LLMs (i.e., input), where LLMs can read and write through actions like reasoning [963]. While long-term memory can be mapped to the external storage like vector databases [539], where LLMs can read through retrieval and write with reflection [688]. Specially, profiles are usually implemented with long-term memory, which is an important feature for an agent that specifies its role and function [821]. The planning component is responsible for generating the action plan based on the information from the memory component. In data format, the plan usually takes the form of text-based instructions [434] or code-based programs [436]. To generate it, LLM-based agents will first propose several candidates and then select a more suitable one among them [429]. The initial plan can be further refined with execution feedback from the environment [530]. The execution component is in charge of carrying out the plan from the planning component, which can be fulfilled by the internal LLM [434] or external tools [963].